OpenStudy (anonymous):

The Natural Environmental Agency is looking for someone who can design a new retaining pond for an endangered species of toucans. The pond needs to be rectangular in shape, and must be 48 square miles in size. The Agency wants the length of the pond to be 8 miles less than its width, but doesn’t know what that means the specific side lengths of the rectangle need to be

OpenStudy (anonymous):

4 and 12

OpenStudy (anonymous):

so 4 and 12 is the answer how did u get it?

OpenStudy (anonymous):

so just suppose a side to be x then second one will be x+8 if u get this then i will go ahead to explain next step

OpenStudy (anonymous):

so 8x

OpenStudy (anonymous):

no x+8

OpenStudy (anonymous):

i dont get it :(

OpenStudy (anonymous):

The Agency wants the length of the pond to be 8 miles less than its width

OpenStudy (anonymous):

so if u suppose the length to be x then width will be x+8

OpenStudy (anonymous):

did u get it now or i should try in another way

OpenStudy (anonymous):

i get it now

OpenStudy (anonymous):

ok then u need to find the area of the rectangle area of rectangle= length*width so area= x * (x+8)

OpenStudy (anonymous):

and thats how you get 4 & 12

OpenStudy (anonymous):

and the area is equal to 48 so area=x * (x+8) =48 x^2+8x=48 x^2+8x-48=0

OpenStudy (anonymous):

thanks !

OpenStudy (anonymous):

can u solve after that.....!!

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

ok try

OpenStudy (anonymous):

if u cant let me know...!!

OpenStudy (anonymous):

X1 = 4 X2 = -12

OpenStudy (anonymous):

is that correct

OpenStudy (anonymous):

{4,-12}

OpenStudy (anonymous):

yeah that is correct so here u will take the value of x=4 for ur solution in the question u have to drop x=-12 because the value of a rectangle can not be negative so -12 is of no use

OpenStudy (anonymous):

try to understand what i m saying...

OpenStudy (anonymous):

so we turn it into a positive or?

OpenStudy (anonymous):

no just understand by solving the equation we have found 2 values of x that is 4 and -12 here in our question side of a rectangle cant be negative so we will drop -12 so the length of the rectangle will be x=4 when u understand this i will tell u how to find width

OpenStudy (anonymous):

ok i get it now u can only hav 1 value for x

OpenStudy (anonymous):

yes now u need to recall what lengths we supposed for our rectangles just to solve it if u suppose the length to be x then width will be x+8

OpenStudy (anonymous):

which x=4 so 8+4 =12

OpenStudy (anonymous):

yes that is correct u cant change sign according to ur need there must be a procedure for everything u doo thats totally goood.....i hope i helped....!! :)